Compound amount


what is the compound amount and compound interest of 360,000 invested at 12% compounded semi-annually for 7 years


1
Expert's answer
2021-03-17T10:59:36-0400

Compound amount: "A=360000(1+\frac{0.12}{2})^{7*2}=813925.42".

Compound interest: "I=813925.42-360000.00=453925.00."